# DynEx: Dynamic Code Synthesis with Structured Design Exploration for Accelerated Exploratory Programming

### v1

@@ Abstract.
Recent advancements in large language models have significantly expedited the process of generating front-end code. This allows users to rapidly prototype user interfaces and ideate through code, a process known as exploratory programming. However, existing LLM code-generation tools focus more on technical implementation details rather than finding the right design given a particular problem. We present DynEx, an LLM-based method for design exploration in accelerated exploratory programming. DynEx uses LLMs to guide users through a structured Design Matrix to explore the design space before dynamic iterative implementation. It also introduces a technique to self-invoke generative AI, enabling the creation of a diverse suite of applications. A user study of 10 experts found that DynEx increased design exploration and enabled the creation of more complex and varied prototypes compared to a Claude Artifact baseline. We conclude with a discussion of the implications of design exploration for exploratory programming.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing User interface programming
Figure 1. DynEx is an LLM-based system for exploratory programming. It guides users through a design space using a structured Design Matrix. Based on the Design Matrix content, DynEx brainstorms the necessary requirements needed to build the UI, generates a detailed project spec to translate the abstract concepts into an implementation plan, and generates synthetic placeholder data to better prototype the application. DynEx then breaks down the project into tasks, which users can iteratively implement until they create a working prototype. Users can incrementally add features until they are satisfied, and also create a variety of new prototypes using the system.
@@ 1. Introduction
Recently released large language models (LLMs) have shown remarkable capabilities in generating code, particularly front-end code  (Si et al., 2024). This allows users to get ideas off the ground faster by building code-based user-interface (UI) prototypes and testing them directly in real world contexts, a process known as exploratory programming. Exploratory programming is crucial for experimental projects where real design testing is more productive than upfront specifications. Many interactions are difficult to simulate in low-fidelity prototyping mediums, particularly when testing data-driven applications (Beth Kery and Myers, 2017); prototyping is crucial in these cases to truly test ideas. LLM code-generation abilities present a unique opportunity to accelerate exploratory programming and ideate through code.
@@
Existing tools that leverage LLM code-generation capabilities, such as GPTPilot  (Pythagora.io, 2024) and OpenHands (formerly OpenDevin)  (Wang et al., 2024), predominantly focus on code development rather than creating an application that considers the end-user experience. GPTPilot prompts the user to consider implementation-level details, such as which packages to use, rather than broader aspects of the user’s ideas, like who the application is for, what problem the application is trying to solve, and what the core methodology is for guiding the solution. Claude Artifact is more suitable for exploration as it enables users to chat with the LLM, but its fundamentally unstructured approach can also be limiting. Users easily become fixated on implementation over exploration, appending features onto prototypes linearly rather than using the chat-bot to laterally exploring their problem space  (Anthropic, 2024). Ultimately, emphasis on implementation over user-centric design can lead to design fixation  (Alipour et al., 2017); users may be led down one path of technical implementation, but lack the divergent exploration that can enable them to find the most fitting approach for their problem  (C.S. Durão et al., 2018). These applications may function well technically, but can lack the depth of refinement that helps ground a prototype in real user needs.
@@
We present DynEx, an LLM-based method for exploratory programming for functional UIs. DynEx helps the user build code-based UI prototypes in two stages: (1) structured design exploration through a Design Matrix, and (2) dynamic iterative implementation with LLM code synthesis. The Design Matrix brainstorms unique ideas through through idea generation to help users explore the design space. The idea is then grounded, a process to synthesize abstract ideas into detailed application designs. The system next breaks down the project into steps to execute iteratively, and generates code that self-invokes multi-modal LLMs to produce applications that have larger sets of synthetic placeholder data, can generate images, and provide recommendations, enabling the creation of a diverse suite of applications.
@@
  • DynEx, an LLM system for accelerated exploratory programming where the user inputs a problem they want to solve and rapidly creates UI prototypes. The system guides users through a design space, then allows users to iteratively generate and interact with UI code. Users can create multiple variations of prototypes and iteratively refine their ideas.
@@
  • Self-invoked multi-modal LLMs, a technique for building code-based UI prototypes where the generated code can call other LLMs to create a diverse and rich suite of applications.
  • A qualitative evaluation of 10 experts, demonstrating that DynEx enabled them to explore the design space and create more complex and varied prototypes as compared to a Claude Artifact baseline where users must lead the design process themselves.
@@ 2.1. Exploratory Programming and Prototyping
Exploratory programming is a practice in which programmers actively experiment with different possibilities using code  (Beth Kery and Myers, 2017). It is grounded by 5 main characteristics: (1) Needs for Exploration, (2) Code Quality Tradeoffs, (3) Ease or Difficulty of Exploration, (4) Exploration Process, and (5) Group or Individual Exploration. Like other prototyping methods, it is important to consider the cost of exploratory programming compared to its value in providing insights  (Tiong et al., 2019). Exploratory programming for UI creation must be as easy and quick as possible while still creating functional code. It must also produce sufficiently complex interfaces that can actually inform designers on how users will interact with them. Finally, it is crucial that exploratory programming processes support lightweight version control so that individuals can easily see the variations they create  (Kery et al., 2017).
@@
Prototyping and exploratory programming are forms of creative problem-solving. During exploratory programming, designers iteratively test and refine ideas via code. The ability to generate multiple alternatives, explore directions, and revert to previous versions have been labeled as requirements for creativity support tools (Shneiderman, 2007). Past research demonstrates the effectiveness of prototyping during design (Dow et al., 2011; C.S. Durão et al., 2018). Learning theory research suggests that generating variations of outputs enables critical reflection and deepens understanding in problem domains (Marton and Booth, 1997). Prototyping has been demonstrated to be effective in serving as a means of inquiry to increase granularity of design requirements based on user reactions  (Wensveen and Matthews, 2014; Kang et al., 2023). However, it is not always possible to prototype UIs in low-fidelity mediums. Applications often change requirements and need to support complex user interactions  (Chen and Yoon, 2024). Even higher-fidelity approaches like Figma  (Figma, Inc., 2023) also have limitations in modeling the interactive experience of a UI. It can lack the richness of user experience that comes from even simple backend computation or data processing  (Beth Kery and Myers, 2017) and have other time and resource costs  (Tiong et al., 2019). Exploratory programming addresses these issues, allowing for the prototyping of code-based interactive applications.
@@ 2.2.1. LLM-Based Code Synthesis Tools
State of the art models for code generation include GPT-4, AlphaCode, CodeGEN, Code Llama, and Gemini  (OpenAI, 2023; Li et al., 2022; Nijkamp et al., 2022; Meta AI, 2023; Google, 2023). These models generally perform well in transforming a natural language problem specification into a simple code solution. They are generally most effective at creating code segments for specific functions or features. Systems built around these LLMs show promise for improving LLM-code generation beyond direct prompting. However, many of the UI code generation tools have limitations. For example, Co-Pilot  (GitHub, 2024) is a system that can modify existing code repositories, but it does not support end-to-end code generation. OpenHands  (Wang et al., 2024) and GPTPilot  (Pythagora.io, 2024) are open-source agent driven systems for end-to-end software development. They use separate LLM agents to represent various roles within the software development process and implement a technical specification. However, these systems lack the support for exploratory programming that is embedded in DynEx, which helps users explore during the planning and problem specification stages of development.
@@ 2.2.2. LLM-Based Code Synthesis Tools for UIs
LLMs have been proven to be especially good at creating UI code. Specifically, LLMs have been shown to be extremely apt at writing code from provided UI screenshots and designs  (Lu et al., 2023; Soselia et al., 2023; Si et al., 2024; Xiao et al., 2024; Qian et al., 2024; Wu et al., 2024; Wan et al., 2024; Yun et al., 2024). LLMs have also shown the ability to write UI code from natural language prompts; WebSim is a system which enables users to created simulated UIs with minimal prompting  (WebSim.ai, 2023). However, these systems do not offer users much design support, placing the brunt of ideation and design space exploration on the user. DynaVis is a tool that dynamically synthesizes widgets for data visualization UI. While it allows users to make dynamic components, it does not support the creation of a complete user interface (Vaithilingam et al., 2024).
@@
Claude Artifact  (Anthropic, 2024) likely offers the best out-of-the-box solution for UI prototyping. It creates functional UIs from conversational prompts, provides users with a window for rendering, and keeps track of version history. However, Claude Artifact does not guide users through the design exploration process. The onus is on the user to use repeated prompting to add new features and move towards a solution. Given that LLMs have demonstrated the ability to effectively power creativity support tools and perform divergent design space exploration in other domains  (Suh et al., 2024; Yuan et al., 2022; Petridis et al., 2023), it should be possible to build LLM systems that support both design exploration and implementation for UI prototyping.
2.3. Informing the Development of LLM-Based Creativity Support Tools
It is necessary for LLM-Based creativity support tools to have sufficient structures for users to create and compare variations within the design space  (Shneiderman, 2007). Moreover, users need to be able to think through all dimensions of their problem, which is a challenging task for people to do on their own  (MacNeil et al., 2017). When left to their own devices, people tend to prematurely converge on ideas before thinking through all possibilities  (Cross, 2004; Dow et al., 2011; Janis, 1982; Jansson and Smith, 1991).
Prior work enables thorough design space exploration by helping users define dimensions to explore. Luminate  (Suh et al., 2024) is a system for exploring design spaces for narratives based on dimensions that can characterize the design space. Users can arrange and cluster generated outputs based on dimensions such as genre, tone, or setting to structure the design space. Other LLM-based writing support tools use pre-defined dimensions such as feasibility and usability, as opposed to synthesizing dimensions on the fly  (Gmeiner and Yildirim, 2023). We realize this dimensional thinking in DynEx’s Design Matrix. By having users traverse through the Design Matrix before any implementation occurs, DynEx forces users to thoroughly think through all of the dimensions of the problem.
@@ 2.3. Informing the Development of LLM-Based Creativity Support Tools
Self-referential multi-modal LLMs have been demonstrated to improve the outputs of LLM-based systems. WebSim  (WebSim.ai, 2023) utilizes self-referential LLMs to generate images for created UIs. Jigsaw is a system for prototyping that utilizes puzzle pieces as metaphors to represent AI foundational models  (Lin and Martelaro, 2024). This representation allows users to understand and chain foundational model capabilities for the creation of more complex outputs. We build upon self-referential LLMs in DynEx by self-invoking methods that can allow for the creation of images or placeholder data, and emulate recommendation or sensemaking systems to power the UI prototypes.
There are several other factors we consider more broadly that are relevant to LLM-systems. First, many tasks are too complex for LLMs to accomplish in one prompt and have to be broken down into smaller problems  (Wu et al., 2022, 2021). LLMs are also more successful in accomplishing tasks when asked to create a plan and reason through intermediate steps  (Cai et al., 2024; Kojima et al., 2023; Suzgun et al., 2022; Wei et al., 2022). Creating a complete UI is clearly a involved task; DynEx breaks down the project into steps to create an implementation plan that can be iteratively executed. Second, tasking non-AI experts with writing many trivial prompts can reduce the effectiveness of LLM-systems  (Cai et al., 2024; Zamfirescu-Pereira et al., 2023). In prior prototyping studies, users spent excessive time debugging and iterating on prompts  (Jiang et al., 2022); thus, DynEx has structured prompting that enables users to focus their energy on expressing their ideas as opposed to prompt engineering for functional outputs. Finally, the difficulty in sourcing example data has limited the success in prototyping with LLMs. (Jiang et al., 2022). DynEx therefore has specific functionality to self-invoke LLMs and easily create synthetic data for applications.
3. System
@@
We present DynEx, a LLM-based system for accelerated exploratory programming that assists users with design exploration and code implementation. The input is the problem that the user wants to solve. It can be vague idea, such as an application to help with plant-watering or meal-planning; something that the user does not have a clear solution for. The output is a functional UI that can support placeholder data, images, and dynamic data. It is an HTML page written in Javascript, HTML, and CSS, that also utilizes React  (Meta Platforms, Inc., 2013) and Material UI Library  (MUI Core Contributors, 2014), a comprehensive React component library offering foundational elements like buttons, dropdowns, input forms, and more.
@@
DynEx is designed for programmers who have personal projects and unrealized ideas they want to explore. Many of these ideas remain underdeveloped or unrealized for years, if not indefinitely. Bringing these ideas to fruition is challenging because it not only requires developer experience to prototype the idea in code, but also requires addressing gaps in the design. DynEx helps concretize users’ ideas through design exploration and implementation, allowing them to take a crucial step towards creating a minimal viable product (MVP).
The pipeline has 2 steps:
@@
    Design Exploration through the Design Matrix, where the system helps the user identify the person, approach, and interaction paradigm of the application using idea generation and grounding, a method of specifying ideas to become a more solidified design.
@@
    Dynamic Iterative Implementation, where the system generates code iteratively, allowing for user input at every step, until the creation of the final prototype. Our implementation process supports the generation of synthetic placeholder data and self-invokes multi-modal LLMs, a technique for the generated code to call other LLMs to create a dynamic and diverse suite of applications.
The system was written in Python, Typescript, and Flask. We use GPT-4 in our Design Matrix for design exploration – due to its vast understanding across various domains, it is well suited for brainstorming and refining ideas. We use Claude 3.5 Sonnet for code generation – Claude is known to be better at writing UI code, and it enables us to use Claude Artifact as a baseline for comparison in our evaluation. Claude is also highly responsive to explicit instructions, which is crucial for code-generation.

### v3

@@ Abstract.
Recent advancements in large language models have significantly expedited the process of generating front-end code. This allows users to rapidly prototype user interfaces and ideate through code, a process known as exploratory programming. However, existing LLM code generation tools focus more on technical implementation details rather than finding the right design given a particular problem. We present DynEx, an LLM-based method for design exploration in accelerated exploratory programming. DynEx introduces a technique to explore the design space through a structured Design Matrix before creating the prototype with a modular, stepwise approach to LLM code generation. Code is generated sequentially, and users can test and approve each step before moving onto the next. A user study of 10 experts found that DynEx increased design exploration and enabled the creation of more complex and varied prototypes compared to a Claude Artifact baseline. We conclude with a discussion of the implications of design exploration for exploratory programming.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714115isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing User interface programming
Figure 1. DynEx is an LLM-based system for exploratory programming. It guides users through a design space using a structured Design Matrix. Based on the Design Matrix content, DynEx brainstorms the necessary requirements needed to build the UI, generates a detailed project specification to translate the abstract concepts into an implementation plan, and generates data to better prototype the application. DynEx then breaks down the project into tasks, which users can iteratively implement until they create a working prototype. Users can incrementally add features until they are satisfied, and also create a variety of new prototypes using the system.
@@ 1. Introduction
Recently released large language models (LLMs) have shown remarkable capabilities in generating code, particularly front-end code  (Si et al., 2024). This allows users to get ideas off the ground faster by building code-based user-interface (UI) prototypes and testing them directly in real world contexts, a process known as exploratory programming. Exploratory programming is crucial for experimental projects where real design testing is more productive than upfront specifications. Many interactions are difficult to simulate in low-fidelity prototyping mediums, particularly when testing data-driven applications (Beth Kery and Myers, 2017). Prototyping is crucial in these cases to truly test ideas. LLM code-generation abilities present a unique opportunity to accelerate exploratory programming and ideate through code.
@@
Existing tools that leverage LLM code-generation capabilities, such as GPTPilot  (Pythagora.io, 2024) and OpenHands (formerly OpenDevin)  (Wang et al., 2024b), predominantly focus on code development rather than creating an application that considers the end-user experience. GPTPilot prompts the user to consider implementation-level details, such as which packages to use, rather than broader aspects of the user’s ideas: who the application is for, what problem the application is trying to solve, and what the core methodology is for guiding the solution. Claude Artifact is more suitable for exploration as it enables users to chat with the LLM, but its fundamentally unstructured approach can also be limiting. Users easily become fixated on implementation over exploration, appending features onto prototypes linearly rather than using the chatbot to laterally exploring their problem space  (Anthropic, 2024). Ultimately, emphasis on implementation over user-centric design can lead to design fixation  (Alipour et al., 2017); users may be led down one path of technical implementation, but lack the divergent exploration that can enable them to find the most fitting approach for their problem  (C.S. Durão et al., 2018). These applications may function well technically, but can lack the depth of refinement that helps ground a prototype in real user needs.
@@
We present DynEx, an LLM-based method for exploratory programming for functional UIs. DynEx helps the user build code-based UI prototypes in two stages: (1) structured design exploration through a Design Matrix, and (2) LLM code generation through Modular Stepwise Implementation. The Design Matrix brainstorms unique ideas across three dimensions (Person, Approach, and Interaction) to help users explore the design space. The idea is then grounded, a process to synthesize abstract ideas into detailed application designs. The system next breaks down the project into modular steps to execute sequentially. Each step is distinct and nonoverlapping, allowing users to modify and test code at each step without worrying about unintended changes affecting other steps. Our prototypes self-invokes multi-modal LLMs to produce applications that have larger sets of synthetic placeholder data, can generate images, and provide recommendations, enabling the creation of a diverse suite of applications.
@@
  • DynEx, an LLM system for accelerated exploratory programming where the user inputs a problem they want to solve and rapidly creates UI prototypes. The system guides users through a design space, then allows users to sequentially generate and interact with UI code. Users can create multiple variations of prototypes and iteratively refine their ideas.
@@
  • Modular Stepwise Implementation, a simple technique for building code-based UI prototypes where the implementation plan is broken down into steps and code is generated sequentially. At each step, users can edit or approve the code before moving onto the next step. Each step is modular, meaning that it is distinct, nonoverlapping, and testable.
  • A qualitative evaluation of 10 experts, demonstrating that DynEx enabled them to explore the design space and create more complex and varied prototypes compared to a Claude Artifact baseline where users must lead the design process themselves.
@@ 2.1. Exploratory Programming and Prototyping
Exploratory programming is a practice in which programmers actively experiment with different possibilities using code  (Beth Kery and Myers, 2017). It is grounded by five main characteristics: (1) Needs for Exploration, (2) Code Quality Tradeoffs, (3) Ease or Difficulty of Exploration, (4) Exploration Process, and (5) Group or Individual Exploration. Like other prototyping methods, it is important to consider the cost of exploratory programming compared to its value in providing insights  (Tiong et al., 2019). Exploratory programming for UI creation must be easy and quick but still produce functional code. It must also produce sufficiently complex interfaces that can actually inform designers on how users will interact with them. Finally, it must support lightweight version control so that individuals can easily see the variations they create  (Kery et al., 2017).
@@
Prototyping and exploratory programming are forms of creative problem-solving. During exploratory programming, designers iteratively test and refine ideas via code. Generating multiple alternatives, exploring directions, and reverting to previous versions have been labeled as requirements for creativity support tools (Shneiderman, 2007). Past research demonstrates the effectiveness of prototyping during design (Dow et al., 2011; C.S. Durão et al., 2018). Learning theory research suggests that generating variations of the output enables critical reflection and deepens understanding in problem domains (Marton and Booth, 1997). Prototyping has been shown to be an effective means of inquiry to increase the granularity of design requirements based on early user feedback  (Wensveen and Matthews, 2014; Kang et al., 2023). However, it is not always possible to prototype UIs in low-fidelity mediums. Applications often change requirements and need to support complex user interactions  (Chen and Yoon, 2024). Even higher-fidelity approaches like Figma  (Figma, Inc., 2023) have limitations in modeling the interactive experience of a UI. It can lack the richness of user experience that comes from simple backend computation or data processing  (Beth Kery and Myers, 2017) and have other time and resource costs  (Tiong et al., 2019). Exploratory programming addresses these issues, allowing for the prototyping of code-based interactive applications.
2.2. Design Space Exploration with AI
HCI literature describes four classic classes of design space exploration (Shireen, 2020; Davis et al., 2024). First, example-based exploration helps users ideate solutions by searching a database of related examples to inspire them (Chou et al., 2023; Koch et al., 2019; Kang et al., 2021) or help find flaws in their ideas or applications (Wang et al., 2024a). Second, reflective exploration provides a rich history-keeping mechanism that facilitates reflection during ideation when needed (Klemmer et al., 2002). Third, genetic exploration generates new solutions by combining components of existing examples (Kim et al., 2022). Finally, dimensional exploration provides a mechanism for breaking down a design space into fundamental dimensions. It first explores multiple ideas for each dimension separately, then synthesizes ideas from each dimension to form a cohesive solution (Micheli et al., 1997; Walker and Thomas, 1997).
Dimensional design exploration has been investigated across various systems. It poses a significant challenge for individuals attempting it independently  (MacNeil et al., 2017). When left to their own devices, people tend to prematurely converge on ideas without fully considering all possibilities  (Cross, 2004; Dow et al., 2011; Janis, 1982; Jansson and Smith, 1991). Existing systems use LLMs to help with dimensional design space exploration. For example, Spellburst (Angert et al., 2023) helps users navigate a multi-dimensional space of generative art possibilities for refining the current design. Luminate (Suh et al., 2024a) facilitates the initial exploration of narrative design spaces for users by generating dimensions with LLMs; however, the dimensions are not controlled and sometimes result in incoherent outputs. We focus on the dimensional approach for initial design space exploration through DynEx’s Design Matrix, which allows users to thoroughly think through their problem in three fixed dimensions (Person, Approach, Interaction) before implementation, which is showcased in Section 3. The already-defined dimensions in the Design Matrix are used as context to brainstorm relevant ideas for the current dimension, so that the separate dimensions synthesize into a coherent prototype design.
@@ 2.3.1. LLM-Based Code Synthesis Tools
State of the art models for code generation include GPT-4, AlphaCode, CodeGEN, Code Llama, and Gemini  (OpenAI, 2023; Li et al., 2022; Nijkamp et al., 2022; Meta AI, 2023; Google, 2023). These models generally perform well in transforming a natural language problem specification into a simple code solution. They are most effective at creating code segments for specific functions or features. Systems built around these LLMs show promise for improving LLM-code generation beyond direct prompting. However, many of the UI code generation tools have limitations. For example, Co-Pilot  (GitHub, 2024) is a system that can modify existing code repositories, but it does not support end-to-end code generation. OpenHands  (Wang et al., 2024b) and GPTPilot  (Pythagora.io, 2024) are open-source agent-driven systems for end-to-end software development. They use separate LLM agents to represent various roles within the software development process and implement a technical specification.
@@ 2.3.2. LLM-Based Code Synthesis Tools for UIs
LLMs have been proven to be especially good at creating UI code. Specifically, LLMs have been shown to be extremely apt at writing code from provided UI screenshots and designs  (Lu et al., 2023; Soselia et al., 2023; Si et al., 2024; Xiao et al., 2024; Qian et al., 2024; Wu et al., 2024; Wan et al., 2024; Yun et al., 2024) and natural language prompts; WebSim is a system which enables users to created simulated UIs with minimal prompting  (WebSim.ai, 2023). However, these systems do not offer users much design support, placing the brunt of ideation and design space exploration on the user. DynaVis is a tool that dynamically synthesizes widgets for data visualization UI. While it allows users to make dynamic components, it does not support the creation of a complete user interface (Vaithilingam et al., 2024).
@@
Claude Artifact  (Anthropic, 2024) likely offers the best out-of-the-box solution for UI prototyping. It creates functional UIs from conversational prompts, provides users with a window for rendering, and keeps track of version history. However, Claude Artifact does not guide users through the design exploration process. The onus is on the user to use repeated prompting to add new features and move towards a solution. Given that LLMs have demonstrated the ability to effectively power creativity support tools and perform divergent design space exploration in other domains  (Suh et al., 2024b; Yuan et al., 2022; Petridis et al., 2023), it should be possible to build LLM systems that support both design exploration and implementation for UI prototyping.
2.3.3. Tool-Calling with AI
@@
Tool-calling has been demonstrated to improve the outputs of LLM-based systems. It refers to the ability to interact with external tools, APIs, plugins, or even other LLMs to enhance their functionality beyond text-based responses (Yang et al., 2023). The difficulty in sourcing example data has limited the success in prototyping with LLMs  (Jiang et al., 2022). Tool-calling is integral in making LLMs more interactive and useful in complex, real-world applications. WebSim calls LLMs to generate images for created UIs  (WebSim.ai, 2023). Jigsaw is a system for prototyping that uses puzzle pieces as metaphors to represent AI foundational models  (Lin and Martelaro, 2024), allowing users to chain foundational model capabilities for the creation of more complex outputs. We use tool-calling in DynEx by enabling our prototypes to self-invoke generative AI, allowing for the creation of images, text, or data, and emulate recommendation or sensemaking systems.
2.3.4. Task-Breakdown and Structured Prompting with AI
Many tasks are too complex for LLMs to accomplish in one prompt and have to be broken down into smaller problems  (Wu et al., 2022, 2021). LLMs are also more successful in accomplishing tasks when asked to create a plan and reason through intermediate steps  (Cai et al., 2024; Kojima et al., 2023; Suzgun et al., 2022; Wei et al., 2022). Creating a complete application is a involved task; DynEx decomposes the project into steps to create an implementation plan that can be iteratively executed.
3. System Overview
@@
We introduce DynEx, an LLM-based system for accelerated exploratory programming that assists users with design exploration and code implementation. The input is the problem that the user wants to solve. It can be vague, such as an application to help with plant-watering or meal-planning. The output is a functional React application that addresses the user’s problem. It integrates generative AI to create images and data that the application requires, enabling the creation of a realistic prototype suitable for user interaction. The code is contained within a single file HTML page written in Javascript, HTML, and CSS, and is limited to 450 lines of code due to the 4096 token constraint from Claude 3.5 Sonnet, which we used for code generation.
@@
DynEx is designed for programmers who have personal projects they want to explore. Many of these ideas remain underdeveloped or unrealized for years, if not indefinitely. Bringing these ideas to fruition is challenging because it requires not only developer experience to prototype the idea in code, but also addressing gaps in the design. DynEx helps concretize users’ ideas through design exploration and implementation, allowing them to take a crucial step towards creating a minimal viable product (MVP).
The system has 2 stages:
@@
    Design Exploration through the Design Matrix, where the system helps the user identify the person, approach, and interaction paradigm of the application.
@@
    Code Generation through Modular Stepwise Implementation, where the system creates a multi-step plan with modular steps, and code is generated for each step sequentially. The user tests and approves the code at each step before the system generates code for the next step.
This remainder of this section introduces the main concepts behind the Design Matrix and Modular Stepwise Implementation. Section 4 then illustrates a usage scenario featuring a user developing a prototype with DynEx, followed by Section 5, which discusses DynEx’s implementation details.